# Electrical networks and data analysis in phylogenetics

### v1

@@ Abstract.
It has been known for a while that the resistance matrix defines a metric on the nodes of the electrical networks. Moreover for a circular electrical network this metric obeys the Kalmanson property as it was shown recently. We will call such a metric an electrical Kalmanson metric. The main results of this paper is a complete description of the electrical Kalmanson metrics in the set of all Kalmanson metrics in terms of the geometry of the positive Isotropic Grassmannian whose connection to the theory of electrical networks was discovered earlier.
@@ 1. Introduction
In the case where there are more than two nodes, there is an important generalization of $R_{ij}$. Given an electrical network with $n$ nodes, there is a linear endomorphism of the vector space of functions defined on the nodes, constructed as follows: for each such a function, there is a unique extension off to all the vertices which satisfies Kirchhoff’s current law at each interior vertex. This function then gives the current $I$ in the network at the boundary vertices defining a linear map which is called the Dirichlet-to-Neumann map or the network response. The matrix of this map is called the response matrix. It plays a key role in the theory and applications of the electrical network [9].
@@
A classic problem in data analysis is studying the systems of subsets defined by either a similarity or a dissimilarity function on $X$ which is either observed directly or derived from a data set. While the latter makes use of splits and split metrics, the key ingredients of the former are systems of clusters, subsets of $X$, and elementary similarity functions. One can interpret splits as distinctive features and clusters as common features, see [1], [2] for an introduction to these ideas.
We argue that the response and the resistance matrices should be viewed as a similarity and a dissimilarity function on the set of nodes of an electrical network and as such they are related via the covariance mapping also known as the Farris transform or the Gromov product.
@@
In computational biology one can construct a distance between two species (such as Hamming distance) which records the proportion of characters where the two species differ. Such a record can be encoded by a real symmetric, nonnegative matrix called a dissimilarity matrix. An important problem in phylogenetics is to reconstruct a weighted tree which represents this matrix. In most cases, a tree structure is too constraining. The notion of a split network is a generalization of a tree in which multiple parallel edges signify divergence. A geometric space of such networks was introduced [14], forming a natural extension of the work by Billera, Holmes, and Vogtmann on tree space. It has been studied from different points of view since it is related to a number of objects in mathematics: the compactification of the real moduli space of curves, to the Balanced Minimal Evolution polytopes and Symmetric Traveling Salesman polytopes to name a few.
@@
The appropriate metric on the set of species defined by a split network has a very special property called the Kalmanson property which distinguishes it completely in the set of all metrics [1].
It has been discovered recently that the resistance metric defined by a circular planar electric network obeys the Kalmanson property [15]. We will call such a split metric an electrical Kalmanson metric. The main results of this paper is a complete description of the set of the electrical Kalmanson metrics inside the set of the Kalmanson metrics. For this we exploit the connection between the space of electrical networks and the positive Isotropic Grassmannian $\mathrm{IG}_{\geq 0}(n-1,2n)$ investigated in [24], [5], [4], [7]. It turns out that the Kalmanson property itself is a consequence of this connection and the description of the electrical Kalmanson metrics we provide is given entirely in terms of geometry of the positive part of this projective variety. It is remarkable that the theory of positivity might play a role in studying phylogenetic networks since it would allow us to apply the powerful machinery of the theory of cluster algebras developed for describing positivity in mathematical objects [27].
@@
All this story should be extended to the compactifications of the respective spaces, taking cactus networks, the known strata in the compactification of electrical networks, to the newly defined compactified split systems as it was started in [13]. In this picture the cactus networks should correspond to the pseudometrics playing the role of dissimilarity functions.
@@
The authors are grateful to Borya Shapiro, Misha Shapiro, Anton Petrunin and Lazar Guterman for useful disscussions and suggestions.
@@ 2.1. The space of electrical networks
A connected electrical network $\mathcal{E}$ is a connected graph with positive weights (conductances) attached to the edges and a chosen subset of the set of vertices which are called the boundary vertices or the nodes.
@@
There are many electrical networks which have the same response matrix, we will describe them now. The following five local network transformations given below are called the electrical transformations. Two electrical networks are said to be equivalent if they can be obtained from each other by a sequence of the electrical transformations. This is an equivalence relation of course, so the set of electrical networks is split in the equivalences classes.
Theorem 2.2 ([9]).
The electrical transformations preserve the response matrix of an electrical network.
Any two circular electrical networks which have the same response matrix are equivalent.
@@
In this paper we will deal mostly with a particular type of connected electrical networks, the connected circular electrical networks. For these we require in addition the graph of $\mathcal{E}$ to be planar and such that the nodes are located on a circle and enumerated clockwise while the rest of the vertices are situated inside of this circle.
We will denote by $E_{n}$ the set of equivalences classes of the connected circular electrical networks.
The set $E_{n}$ allows the following elegant description.
@@
Let $P=(p_{1},\ldots,p_{k})$ and $Q=(q_{1},\ldots,q_{k})$ be disjoint ordered subsets of the nodes arranged on a circle, then $(P;Q)=(p_{1},\ldots,p_{k};q_{1},\ldots,q_{k})$ is a circular pair if $p_{1},\ldots,p_{k},q_{k},\ldots,q_{1}$ are in circular order around the boundary circle. Let $(P;Q)$ be a circular pair then the determinant of the submatrix $M(P;Q)$ whose rows are labeled by $(p_{1},\ldots,p_{k})$ and the columns labeled by $(q_{1},\ldots,q_{k})$ is called the circular minor associated with a circular pair $(P;Q)$.
@@
  • The kernel of $M$ is generated by the vector $(1,1,…,1)$;
@@ 2.2. Cactus electrical networks
Setting an edge conductance to zero or infinity in $\mathcal{E}$ makes sense. According to the Ohm law it means that we delete or contract this edge. Doing it we either get isolated nodes or some nodes get glued together. Will will consider the resulting network as a network with $n$ nodes, remembering how nodes get identified. Such a network is called a cactus electrical network with $n$ nodes. One can think of it as a collection of ordinary circular electrical networks with the total number of the nodes equal to $n$, glued along some of these nodes. Note, that the graph of such a network is planar, but it does not have to be connected.
@@
The definition of a cactus network was introduced in [24] where it was proved that the set $\overline{E}_{n}$ is a compactification of $E_{n}$ in the appropriate sense.
@@ 2.3. Lam embedding
Recall that the real Grassmannian $\mathrm{Gr}(k,n)$ is a differentiable manifold that parameterizes the set of all $k$-dimensional linear subspaces of a vector space $\mathbb{R}^{n}$. In fact it has a structure of a projective algebraic variety. The Pluecker embedding is an embedding of the Grassmannian $\mathrm{Gr}(k,n)$ into the projectivization of the $k$-th exterior power of the vector space $\mathbb{R}^{n}$
@@
For practical calculations one can view the matrix whose rows are the coordinates of the basis vectors $(w_{1},\cdots,w_{k})$ as a representative of this equivalence class. For any ordered sequence $I$ of $k$ positive integers $1\leq i_{1}<\cdots<i_{k}\leq n$ denote by $\Delta_{I}$ the determinant of a $k\times k$ submatrix of the above matrix with columns labeled by the numbers from $I$. The numbers $\Delta_{I}$ are called the Pluecker coordinates of the point $W$ of $\mathrm{G}r(k,n)$. They are defined up to a common non zero factor.
@@
The totally non-negative Grassmannian $\mathrm{Gr}_{\geq 0}(k,m)$ is the subset of the points of the Grassmannian $\mathrm{Gr}(k,n)$ whose Pluecker coordinates $\Delta_{I}$ have the same sign.
The following theorem of T. Lam [24] is one of the key results about the space of electrical nerworks.
@@
where $H$ is a certain subspace of $\bigwedge^{n-1}\mathbb{R}^{2n}$ of dimension equal the Catalan number $C_{n}$.
Moreover the image of the set $E_{n}$ under this bijection is exactly the set of points with the Pluecker coordinates $\Delta_{24\dots 2n-2}$ and $\Delta_{13\dots 2n-3}$ not equal to zero.
Notice that because $M_{R}(\mathcal{E})$ has the rank equal to $n-1$ by 2.4, the dimension of the row space of $\Omega(\mathcal{E})$ is equal to $n-1$, hence it defines a point in $\mathrm{Gr}(n-1,2n)$. The Pluecker coordinates of the point associated with $\Omega(\mathcal{E})$ can be calculated as the maximal size minors of the matrix $\Omega^{\prime}(\mathcal{E})$ obtained from $\Omega(\mathcal{E})$ by deleting, for example, the first row.
We will recall the construction of the embedding of ${E}_{n}$ obtained in [5], which is induced by the above bijection. Let $\mathcal{E}$ be a circular electrical network with the response matrix $M_{R}(\mathcal{E})=(x_{ij}).$ The following $n\times 2n$ matrix
@@
[6] Let $A=(a_{ij})$ be a matrix with satisfies the first three conditions of Theorem 2.4 and $\Omega(A)$ be a matrix constructed out of according to the formular (2.1). If $\Omega(A)$ defines a point in $Gr_{\geq 0}(n-1,2n)$ and the Pluecker coordinate $\Delta_{13\dots 2n-3}\bigl{(}\Omega(A)\bigr{)}$ is not equal to zero, then there is a connected electrical network $\mathcal{E}\in E_{n}$ such that $A=M_{R}(\mathcal{E}).$
@@
For the network $\mathcal{E}$ in $E_{4}$ on the Figure 6 the matrix $\Omega(\mathcal{E})$ has the following form:
@@
In fact the row space of $\Omega(\mathcal{E})$ is isotropic with respect to a particular symplectic form. This refines the above embedding to a submanifold
@@ 3.1. Resistance metric
Let $\Gamma,\omega$ be a connected graph with a weight function $\omega$ on the edges and $n,\,m\in\Gamma$ be two of its vertices. Consider it as an electrical network $\mathcal{E}_{nm}(\Gamma,\omega)$ on a graph $\Gamma$ with a conductivity function $\omega$ by declaring the vertices $n$ and $m$ to be the nodes while the remaining vertices to be internal vertices.
Apply the voltages $U=(U_{n},U_{m})$ to the nodes of $\mathcal{E}_{nm}(\Gamma,\omega)$ such that they induce the boundary currents $I=(1,-1)$. Then the effective resistance between the nodes $n$ and $m$ is defined as
$R_{nm}=|U_{n}-U_{m}|$
Obviously $R_{nm}=R_{mn}$.
@@
[20] Let $\mathcal{E}$ be a connected electrical network with $n$ nodes, and let the boundary voltages $U=(U_{1},\dots,U_{n})$ be such that
@@
Indeed, if the boundary voltages $U$ are such as in (3.2) we can consider all the vertices apart from $i$ and of $j$ as the inner vertices and then $|U_{i}-U_{j}|$ is precisely as in Definition 3.1. ∎
@@
The proposition 3.3 and the lemma 3.2 show that the resistance and the response matrices define each other therefore the theorem 2.2 would hold if we replace the response matrix with the resistance matrix in its statement.
@@
The formula from Proposition 3.3 is well known in different areas of mathematics. It appeared in the literature under the names the Gromov product, the Farris transform, the Covariance mapping between the Cut and the Covariance cones, see [11] for more information. We are planning to explore the properties of the resistance matrix from these points of view in the future publications.
@@
Hence the set of all $R_{vw}$ define a metric on the nodes of $\Gamma$.
@@
Let $\mathcal{E}_{k_{1}k_{2}k_{3}}$ be a connected electrical network obtained from $\mathcal{E}$ by declaring the vertices $k_{1},k_{2},k_{3}$ to be the boundary nodes, while remaining vertices are declared to be inner and $M_{R}(\mathcal{E}_{k_{1}k_{2}k_{3}})$ be its response matrix, than according to Proposition 3.3 we have that:
@@
therefore to get the statement it is enough to verify that $M^{\prime}_{R}(\mathcal{E}_{k_{1}k_{2}k_{3}})^{-1}_{k_{1}k_{2}}\geq 0.$ Indeed, the matrix $M^{\prime}_{R}(\mathcal{E}_{k_{1}k_{2}k_{3}})$ has the following from:
@@
To describe the properties of the resistance metric associated with a connected circular electrical network we will provide a formula for the embedding of $E_{n}$ into $\mathrm{Gr}_{\geq 0}(n-1,2n)$ described in 2.7 which uses the effective resistances matrix instead of the response matrix.
@@
Notice that the matrix $M(R_{\mathcal{E}})=(m_{ij})$ is symmetric and the sum of the matrix entries in each row is zero, in other words it looks like a response matrix of an electrical network. There is a reason for this as it was discovered by R. Kenyon and D. Wilson.
@@
Since $M(R_{\mathcal{E}})$ is degenerated matrix the Pluecker coordinates of a point $Gr(n-1,2n)$ associated with $\Omega_{R}(\mathcal{E})$ can be calculated as the maximal size minors of the matrix $\Omega^{\prime}_{R}(\mathcal{E})$ obtained from $\Omega_{R}(\mathcal{E})$ by deleting, for example, the last row.
@@
[4], [7] The row space of $\Omega_{R}(\mathcal{E})$ defines the same point in the $Gr_{\geq 0}(n-1,2n)$ as the point $\Omega(\mathcal{E})$ defined by the Lam embedding 2.7. In particular the Pluecker coordinate $\Delta_{24\dots 2n-2}(\Omega_{R}(\mathcal{E}))$ is not 0 if and only if $\mathcal{E}$ is a connected circular electrical network.
@@
where $s$ is the shift operator
@@
Many interesting inequalities involving $R_{ij}$ follow from positivity of the Pluecker coordinates of the point represented by $\Omega_{R}(\mathcal{E})$. For some of them we have found an explicit combinatorial meaning, others are still waiting to be interpreted.
@@
Let $\mathcal{E}$ be a connected circular electrical network and let $i_{1},i_{2},i_{3},i_{4}$ be any four nodes in the circular order as it is shown in Figure 3. Then the Kalmanson inequalities hold:
@@
Let $\mathcal{E}_{i_{1}i_{2}i_{3}i_{4}}$ be an electrical network obtained from $\mathcal{E}$ by declaring the vertices $i_{1},i_{2},i_{3},i_{4}$ the boundary nodes, and the rest of the vertices are internal. Obviously $\mathcal{E}_{i_{1}i_{2}i_{3}i_{4}}$ is a connected circular electrical network in $E_{4}$ and let $\Omega^{\prime}_{R}(\mathcal{E}_{i_{1}i_{2}i_{3}i_{4}})$ be its matrix defined by 3.3. By direct computations we conclude that
@@
Let $X$ be a finite set and $D$ is a metric on it. If there is a circular order on $X$ such that the inequalities from the theorem above hold for any four points of $X$ in this circular order we call this metric the Kalmanson metric.

### v2

@@ Abstract.
It has been known for a while that the resistance matrix defines a metric on the nodes of the electrical networks. Moreover for a circular electrical network this metric obeys the Kalmanson property as it was shown recently. We will call such a metric an electrical Kalmanson metric. The main results of this paper is a new description of the electrical Kalmanson metrics in the set of all Kalmanson metrics in terms of the geometry of the positive Isotropic Grassmannian whose connection to the theory of electrical networks was discovered earlier.
@@ 1. Introduction
In the case where there are more than two nodes, there is an important generalization of $R_{ij}$. Given an electrical network with $n$ nodes, there is a linear endomorphism of the vector space of functions defined on the nodes, constructed as follows: for each such a function, there is a unique extension of that function to all the vertices which satisfies Kirchhoff’s current law at each interior vertex. This function then gives the current $I$ in the network at the boundary vertices defining a linear map which is called the Dirichlet-to-Neumann map or the network response. The matrix of this map is called the response matrix. It plays a key role in the theory and applications of the electrical network [9].
@@
A classic problem in data analysis is studying the systems of subsets defined by either a similarity or a dissimilarity function on $X$ which is either observed directly or derived from a data set. While the latter makes use of splits and split metrics, the key ingredients of the former are systems of clusters, subsets of $X$, and elementary similarity functions. One can interpret splits as distinctive features and clusters as common features, see [11] and [27] for an introduction to these ideas.
We argue that the ”inverse” of the response and the resistance matrices should be viewed as a similarity and a dissimilarity function on the set of nodes of an electrical network and as such they are related via the covariance mapping also known as the Farris transform or the Gromov product, see Remark 3.5 for more details.
@@
In computational biology species are naturally assigned collections of symbols from a given set, called characters, and one can construct a distance between two species (such as Hamming distance) which records the proportion of characters where the two species differ. Such a record can be encoded by a real symmetric, nonnegative matrix called a dissimilarity matrix $M$. An important problem in phylogenetics is to reconstruct a weighted tree $T$ with the set of leaves equal to the set of of species and the matrix of the tree metric defined by $T$ on the set of leaves is equal to $M$, see [27] for a nice introduction to these ideas. In most cases, a tree structure is too constraining. The notion of a split network is a generalization of a tree in which a certain type of cycles is allowed. A geometric space of such networks was introduced [14], forming a natural extension of the work by Billera, Holmes, and Vogtmann on tree space. It has been studied from different points of view since it is related to a number of objects in mathematics: the compactification of the real moduli space of curves, to the Balanced Minimal Evolution polytopes and Symmetric Traveling Salesman polytopes to name a few.
@@
The appropriate metric on the set of species defined by a split network has a very special property called the Kalmanson property which distinguishes it completely in the set of all metrics [2], [27].
Stefan Forcey has discovered recently that the resistance metric defined by a circular planar electric network obeys the Kalmanson property [15]. We will call such a split metric an electrical Kalmanson metric. This important result puts a new light on the theory of electrical networks connecting it with phylogenetics and metric geometry.
The purpose of this paper is twofold: firstly we want to collect in one place the relevant facts from the research areas mentioned above and indicate interesting connections between them, secondly we provide a new description of the set of the electrical Kalmanson metrics inside the set of all Kalmanson metrics given in Theorem 3.19. For the later we will exploit the connection between the space of circular planar electrical networks and the non-negative Isotropic Grassmannian $\mathrm{IG}_{\geq 0}(n-1,2n)$ which uses the effective resistance matrix [4], Theorem $5.6$ as opposed to the response matrix. This description is a modification of the original construction given by Thomas Lam [25] and developed further in [5], [7].
It turns out that the Kalmanson property itself is a consequence of this connection and the description of the electrical Kalmanson metrics we provide is given entirely in terms of the geometry of the non-negative part of this projective variety. In our description checking whether a given Kalmanson metric is an electrical Kalmanson metric amounts to checking that the Plücker coordinates of an explicitly defined point in the above Grassmannian are non-negative.
Moreover, using the results from [25] we propose an algorithm for reconstructing a minimal circular planar electrical network out of a given resistance matrix which might be useful for possible applications in phylogenetics.
It is remarkable that the theory of positivity might play a role in studying phylogenetic networks since it would allow us to apply the powerful machinery of the theory of cluster algebras developed for describing positivity in mathematical objects [30].
@@
This story should be extended to the compactifications of the respective spaces, taking cactus networks, the known strata in the compactification of circular planar electrical networks, to the newly defined compactified split systems as it was already started in [13]. In this picture the cactus networks should correspond to the pseudometrics playing the role of dissimilarity functions.
@@
Finally building of an example from [15] we show in Remark 3.22 that the planarity condition on the network is far from being necessary to guarantee the Kalmanson condition for the resistance metric. It leaves the description of the set of electrical Kalmanson metrics as an interesting question.
@@
The authors are grateful to Borya Shapiro, Misha Shapiro, Anton Petrunin, Lazar Guterman and especially to Stefan Forcey for useful discussions and helpful suggestions. The authors are grateful to the referees for carefully reading the paper and suggesting a number of useful comments to make it better.
@@ 2.1. The space of electrical networks
A electrical network $\mathcal{E}$ is a graph $\Gamma$ with positive $\omega$ weights (conductances) attached to the edges and a chosen subset of the set of vertices which are called the boundary vertices or the nodes.
@@
There are many electrical networks which have the same response matrix, we will describe them now. The following five local network transformations given below are called the electrical transformations. Two electrical networks are said to be equivalent if they can be obtained from each other by a sequence of the electrical transformations. This is an equivalence relation of course, so the set of electrical networks is partitioned in the equivalence classes.
Theorem 2.2.
[9] The electrical transformations preserve the response matrix of an electrical network.
@@
In this paper we will deal with a particular type of connected electrical networks, the connected circular planar electrical networks, unless otherwise stated. For these we require in addition the graph of $\mathcal{E}$ to be planar and such that the nodes are located on a circle and enumerated clockwise while the rest of the vertices are situated inside of this circle.
We will denote by $E_{n}$ the set of equivalence classes of the circular electrical networks in this paper. Note that the notation for this set of equivalence classes used in [9] and in [10] is $\Omega_{n}$. The set $E_{n}$ allows the following elegant description.
@@
Let $P=(p_{1},\ldots,p_{k})$ and $Q=(q_{1},\ldots,q_{k})$ be disjoint ordered subsets of the nodes arranged on a circle, then $(P;Q)=(p_{1},\ldots,p_{k};q_{1},\ldots,q_{k})$ is a circular pair if $p_{1},\ldots,p_{k},q_{k},\ldots,q_{1}$ are in non-overlapping circular order around the boundary circle. Let $(P;Q)$ be a circular pair then the determinant of the submatrix $M(P;Q)$ whose rows are labeled by $(p_{1},\ldots,p_{k})$ and the columns labeled by $(q_{1},\ldots,q_{k})$ is called the circular minor associated with a circular pair $(P;Q)$.
@@
  • The kernel of $M$ is generated by the vector $(1,1,\dots,1)$.
Moreover, any two circular electrical networks which have the same response matrix are equivalent.
@@ 2.2. Cactus electrical networks
Setting an edge conductance to zero or infinity in $\mathcal{E}$ makes sense. According to the Ohm law it means that we delete or contract this edge. Doing it we either get isolated nodes or some nodes get glued together. We will consider the resulting network as a network with $n$ nodes, remembering how nodes get identified. Such a network is called a cactus electrical network with $n$ nodes. One can think of it as a collection of ordinary circular electrical networks with the total number of the nodes equal to $n$, glued along some of these nodes. Note, that the graph of such a network is planar, but it does not have to be connected.
@@
The definition of a cactus network was introduced in [25] where it was proved that the set $\overline{E}_{n}$ is a compactification of $E_{n}$ in the appropriate sense. Note that the space of cactus networks in denoted by $E_{n}$ in [25]. We apologize for this unfortunate clash of notations.
@@ 2.3. Lam embedding
Recall that the real Grassmannian $\mathrm{Gr}(k,n)$ is a differentiable manifold that parameterizes the set of all $k$-dimensional linear subspaces of a vector space $\mathbb{R}^{n}$. In fact it has a structure of a projective algebraic variety. The Plücker embedding is an embedding of the Grassmannian $\mathrm{Gr}(k,n)$ into the projectivization of the $k$-th exterior power of the vector space $\mathbb{R}^{n}$
@@
For practical calculations one can view the matrix whose rows are the coordinates of the basis vectors $(w_{1},\cdots,w_{k})$ as a representative of this equivalence class. For any ordered sequence $I$ of $k$ positive integers $1\leq i_{1}<\cdots<i_{k}\leq n$ denote by $\Delta_{I}$ the determinant of a $k\times k$ submatrix of the above matrix with columns labeled by the numbers from $I$. The numbers $\Delta_{I}$ are called the Plücker coordinates of the point $W$ of $\mathrm{G}r(k,n)$. They are defined up to a common non zero factor.
@@
The totally non-negative Grassmannian $\mathrm{Gr}_{\geq 0}(k,m)$ is the subset of the points of the Grassmannian $\mathrm{Gr}(k,n)$ whose Plücker coordinates $\Delta_{I}$ have the same sign or equal to zero.
The following theorem of T. Lam [25] is one of the key results about the space of electrical networks.
@@
where $H$ is a certain subspace of $\bigwedge^{n-1}\mathbb{R}^{2n}$ of dimension equal to the Catalan number $C_{n}$.
Moreover, the image of the set $E_{n}$ under this bijection is exactly the set of points with the Plücker coordinates $\Delta_{24\dots 2n-2}$ and $\Delta_{13\dots 2n-3}$ not equal to zero.
We will recall the explicit construction of the embedding of ${E}_{n}$ obtained in [5], which is induced by the above bijection. Let $\mathcal{E}$ be a circular electrical network with the response matrix $M_{R}(\mathcal{E})=(x_{ij}).$ The following $n\times 2n$ matrix
@@
Since the sum of the entries in each row of $M_{R}(\mathcal{E})$ is equal to 00 (see Theorem 2.4), we have that $\Omega(\mathcal{E})$ has the rank equal to $n-1$. Therefore, the dimension of the row space of $\Omega(\mathcal{E})$ is equal to $n-1$, hence it defines a point in $\mathrm{Gr}(n-1,2n)$. The Plücker coordinates of the point associated with $\Omega(\mathcal{E})$ can be calculated as the maximal size minors of the matrix $\Omega^{\prime}(\mathcal{E})$ obtained from $\Omega(\mathcal{E})$ by deleting, for example, the first row.
@@
[6] Let $A=(a_{ij})$ be a matrix which satisfies the first three conditions of Theorem 2.4 and $\Omega(A)$ be a matrix constructed according to the formula (2.1). If $\Omega(A)$ defines a point in $Gr_{\geq 0}(n-1,2n)$ and the Plücker coordinate $\Delta_{13\dots 2n-3}\bigl{(}\Omega(A)\bigr{)}$ is not equal to zero, then there is a connected electrical network $\mathcal{E}\in E_{n}$ such that $A=M_{R}(\mathcal{E}).$
@@
For the network $\mathcal{E}\in E_{4}$ seen in the Figure 7 the matrix $\Omega(\mathcal{E})$ has the following form:
@@
In fact, the row space of $\Omega(\mathcal{E})$ is isotropic with respect to a particular symplectic form. This refines the above embedding to a submanifold
@@ 3.1. Resistance metric
Now we assume that an electrical network is just connected.
@@
Let $\Gamma,\omega$ be a connected graph with a weight function $\omega$ on the edges and $i,\,j\in\Gamma$ be two of its vertices. Consider it as an electrical network $\mathcal{E}_{ij}(\Gamma,\omega)$ on a graph $\Gamma$ with a conductivity function $\omega$ by declaring the vertices $i$ and $j$ to be the nodes while the remaining vertices to be internal vertices.
Apply the voltages $U=(U_{i},U_{j})$ to the nodes of $\mathcal{E}_{ij}(\Gamma,\omega)$ such that they induce the boundary currents $I=(1,-1)$. Then the effective resistance between the nodes $i$ and $j$ is defined as
$R_{ij}=|U_{i}-U_{j}|$
Obviously $R_{ij}=R_{ji}$.
This definition can be stated which is due to Gustav Kirchhoff in a beautiful way using the combinatorics of the graph. Two pieces of notation are required: recall the generating function of spanning-trees of $\Gamma$ is defined as follows
$T(\Gamma,\omega)=\sum_{T\in T(\Gamma)}\omega(T),$
where the sum is taken over the set $T(\Gamma)$ of all spanning trees of $\Gamma$ and $\omega(T)$ is defined as the product of the weights of the edges of $T$. For a graph $\Gamma,\omega$ as above and vertices $i,j\in V$, we let $\Gamma/ij$ denote the graph obtained by merging the two vertices $i$ and $j$ together into a single vertex.
Theorem 3.2.
(Kirchhoff’s Formula [20], [29]). Let $(\Gamma,\omega)$ be an electrical network, and let $i,j\in V$. The effective resistance between $i$ and $j$ in $\Gamma$ is
$R_{ij}=\frac{T(\Gamma/ij;\omega)}{T(\Gamma;\omega)}$
@@
[21] Let $\mathcal{E}$ be a connected electrical network with $n$ nodes
$\{1,\dots,i,\dots,j,\dots,n\}$, and let the boundary voltages $U=(U_{1},\dots,U_{n})$ be such that
@@
Indeed, if the boundary voltages $U$ are such as in (3.2) we can consider all the vertices except $i$ and $j$ as the inner vertices, and then $|U_{i}-U_{j}|$ is precisely as in Definition 3.1. ∎
@@
Proposition 3.4 and Lemma 3.2 show that the resistance and the response matrices define each other. Therefore, Theorem 2.2 would hold if we replace the response matrix with the resistance matrix in its statement.
@@
The formula from Proposition 3.4 is well known in different areas of mathematics. It appeared in the literature under the names the Gromov product, the Farris transform, the Covariance mapping between the Cut and the Covariance cones, see [11] for more information. This allows us to view $M^{\prime}_{R}(\mathcal{E})^{-1}$, the ”inverse” of the response matrix, as a similarity function corresponding under the Covariance mapping to a dissimilarity function represented by the resistance metric $R_{\mathcal{E}}$. We are planning to explore the properties of the resistance matrix from these points of view in the future publications.
@@
Hence the set of all $R_{vw}$ defines a metric on the nodes of $\Gamma$.
@@
Let $\mathcal{E}_{k_{1}k_{2}k_{3}}$ be a connected electrical network obtained from $\mathcal{E}$ by declaring the vertices $k_{1},k_{2},k_{3}$ to be the boundary nodes, while remaining vertices are declared to be inner and $M_{R}(\mathcal{E}_{k_{1}k_{2}k_{3}})$ be its response matrix, then according to Proposition 3.4 we have that:
@@
therefore to get the statement it is enough to verify that $M^{\prime}_{R}(\mathcal{E}_{k_{1}k_{2}k_{3}})^{-1}_{k_{1}k_{2}}\geq 0.$ Indeed, the matrix $M^{\prime}_{R}(\mathcal{E}_{k_{1}k_{2}k_{3}})$ has the following form:
@@
To describe the properties of the resistance metric associated with a connected circular electrical network, we will provide a formula for the embedding of $E_{n}$ into $\mathrm{Gr}_{\geq 0}(n-1,2n)$ described in Theorem 2.7 which uses the effective resistances matrix instead of the response matrix.
@@
Notice that the matrix $M(R_{\mathcal{E}})=(m_{ij})$ is symmetric and the sum of the matrix entries in each row is zero. In other words, it looks like a response matrix of an electrical network. There is a reason for this as it was discovered by R. Kenyon and D. Wilson.
@@
Since $M(R_{\mathcal{E}})$ is a degenerate matrix, the Plücker coordinates of a point $Gr(n-1,2n)$ associated with $\Omega_{R}(\mathcal{E})$ can be calculated as the maximal size minors of the matrix $\Omega^{\prime}_{R}(\mathcal{E})$ obtained from $\Omega_{R}(\mathcal{E})$ by deleting, for example, the last row.
@@
[4] The row space of $\Omega_{R}(\mathcal{E})$ defines the same point in the $Gr_{\geq 0}(n-1,2n)$ as the point $\Omega(\mathcal{E})$ defined by the Lam embedding, see Theorem 2.7. In particular, the Plücker coordinate $\Delta_{24\dots 2n-2}(\Omega_{R}(\mathcal{E}))$ is not 0 if and only if $\mathcal{E}$ is a connected circular electrical network.
@@
where $s\in\operatorname{Mat}_{2n\times 2n}$ is the shift operator
@@
Many interesting inequalities involving $R_{ij}$ follow from the positivity of the Plücker coordinates of the point represented by $\Omega_{R}(\mathcal{E})$. For some of them we have found an explicit combinatorial meaning, others are still waiting to be interpreted.
@@
Let $\mathcal{E}$ be a connected circular electrical network and let $i_{1},i_{2},i_{3},i_{4}$ be any four nodes in the circular order as it is shown in the Figure 3. Then the Kalmanson inequalities hold:
@@
Let $\mathcal{E}_{i_{1}i_{2}i_{3}i_{4}}$ be an electrical network obtained from $\mathcal{E}$ by declaring the vertices $i_{1},i_{2},i_{3},i_{4}$ the boundary nodes, and the rest of the vertices are internal. Obviously $\mathcal{E}_{i_{1}i_{2}i_{3}i_{4}}$ is a connected circular electrical network in $E_{4}$ and let $\Omega^{\prime}_{R}(\mathcal{E}_{i_{1}i_{2}i_{3}i_{4}})$ be its matrix defined by the formula 3.3. By direct computations we conclude that
@@
Remark 3.10.
Note that the equations 3.6, 3.7 together are equivalent to the tropical inequality:
$R_{i_{1}i_{3}}+R_{i_{2}i_{4}}\geq\max(R_{i_{2}i_{3}}+R_{i_{1}i_{4}},\,R_{i_{1}i_{2}}+R_{i_{3}i_{4}})$
@@
Let $X$ be a finite set and $D$ is a metric on it. If there is a circular order on $X$ such that the inequalities from the theorem above hold for any four points of $X$ in this circular order, we call this metric the Kalmanson metric.